[Hallc_running] [New Logentry] Hall C RC Update: Monday, December 20

gaskelld at jlab.org gaskelld at jlab.org
Mon Dec 20 08:35:01 EST 2021

Logentry Text:
We completed the aluminum dummy and started the LH2 running at the current setting (SHMS left, Q2=6, W=2.4, low epsilon) at about 4:15 pm yesterday.  Beam has been off since 1 am last night due to issues with 1L05 (1L05 is now back up and there appears to be tune beam in the machine, so beam may be back to the hall relatively soon). We have about 5766 good pion events at this setting - I estimate we need about 2 more shifts once beam returns to get full statistics.

The BCM calibration set for this afternoon is still scheduled, but this may be revisited depending on when beam returns.

The will be an RC meeting at 10 am this morning:

RC meeting bluejeans link:

RC meeting page:


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3966335
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211220/0d8bee51/attachment.html>

More information about the Hallc_running mailing list